# CoVoMix2: Advancing Zero-Shot Dialogue Generation with Fully Non-Autoregressive Flow Matching

### v1

@@ 1 Introduction
The field of speech synthesis has witnessed remarkable progress in recent years, particularly in zero-shot text-to-speech (TTS) systems that can generate high-quality, natural-sounding speech in voices not seen during training [1, 2, 3, 4, 5, 6]. These advancements have enabled applications such as personalized virtual assistants, audiobook narration, and interactive voice response systems. However, extending these capabilities to multi-talker dialogue generation, where multiple speakers engage in natural, dynamic, multi-turn conversations, remains a significant challenge.
@@ 2.3 Conversational Speech Synthesis
Conversational Speech Synthesis (CSS) focuses on generating individual utterances that are contextually appropriate within a dialogue, typically from a single speaker. Unlike full dialogue generation, CSS produces each utterance one-by-one, conditioned on previous dialogue history, rather than synthesizing an entire conversation simultaneously [28, 29, 30].
For instance, GPTTalker [31] models multimodal dialogue context by converting history into discrete tokens processed by a language model to generate expressive, context-aware responses. Sesame [8] adopts two AR transformers: a multimodal transformer backbone processes text and audio tokens, followed by a audio decoder transformer that reconstructs high-quality speech. Although Sesame supports dialogue-style synthesis, it fundamentally generates each speaker’s utterance sequentially given the previous dialogue context, rather than modeling simultaneous multi-speaker interaction.
@@ 3.3 Prompt-Level Random Masking
As shown in Figure 1 and 2, for each training sample, we first find all the available monologue segments for each speaker as the prompt candidates. We then randomly select a prompt segment from the candidates. The chosen prompts are concatenated at the very beginning with the masked training sample using a separator token to construct the prompt sequence . Moreover, in the text stream , we use special tokens and to indicate the range of each prompt, instead of providing prompt’s text transcription.
@@ 3.5 Training Strategy: Curriculum Learning and Data Mixing
To improve robustness and generalization, built on the large scale of monologue dataset, we design the data mixing strategy, using various sources of data during the second stage training. Specifically, we mix data from several sources, including ASR-transcribed podcast dialogues, audiobook-style single-speaker datasets and simulated overlapped dialogues to support overlapping capability. Benefited from this data diversity, CoVoMix2 does not require any human-annotated dialogues to get satisfied and natural results. For further enhancement, we demonstrate in Appendix C.2 that fine-tuning the model with just 20 minutes of clean, human-annotated dialogue can significantly boost performance.
@@ 4.1 Training and Inference Data Preparation
To train CoVoMix2, we curated a diverse dataset comprising both in-the-wild dialogues and clean monologue data. The core training corpus consists of 3,000 hours of English podcast data, which we processed with 4 steps, as shown in Figure 3.
First, we applied a market leading speech enhancement API to remove background noise and music. Second, we implemented automatic speech recognition and speaker diarization. The diarization results were used to assign speaker identities to each utterance with timestamps. Specifically, we used the Deepgram API [34]222https://deepgram.com/ to obtain word-level transcriptions and speaker labels. Third, following [9]333https://github.com/vivian556123/NeurIPS2024-CoVoMix/tree/main, long dialogues were segmented into shorter clips, each involving two or one speakers, to improve data quality and training efficiency. Finally, we simulate dialogue segments using the monologue datasets. By pairing utterances from different speakers and introducing controlled overlap or silence ratios, we generate synthetic two-speaker dialogues.
@@
In order to evaluate the model performance, we design a dialogue test set 444https://github.com/vivian556123/covomix2-dialogue-testset.git, containing 1000 dialogue transcriptions from Dailydialog [37] and the acoustic prompts are from Librispeech-test-clean [38]. We also use samples from this dialogue dataset for subjective evaluation. To ensure that the monologue generation capability is retained, we also designed a monologue test set, with results presented in Appendix B.
@@ 4.3 Baseline and Evaluation Metrics
To comprehensively assess the recognition accuracy and speaker consistency, we adopt the following objective evaluation metrics: Real-Time Factor (RTF), Word Error Rate (WER), Speaker-Aware Word Error Rate (SA-WER) [41], Speaker-Aware Speaker Similarity (SA-SIM) and UTMOS [42]. Specifically, We measure the RTF on a single NVIDIA A100 machine. We utilize market leading automatic speech recognition and diarization tool to transcribe the generated speech, and we calculate the SA-WER for each word and their corresponding speaker identity. The SA-SIM enhances SIM by ensuring that the similarity is computed between embeddings attributed to the speaker identity detected by the diarization model. We utilize WavLM-TDNN [43] to extract the speaker embeddings.
We perform a human evaluation on the generated dialogue examples. We conducted a Comparable Mean Opinion Score (CMOS) experiment to assess user preference in terms of speaker turn handling, interactivity, fluency, and coherence. 15 professional linguistic experts provide judges for all subjective evaluations. They provide a rating to the second audio, which is randomly selected from a pair of audios, in the (-3 to +3) range. Detailed instructions are in Appendix D.
@@ 5.1 Objective and Subjective Evaluation Results
In terms of content accuracy and speaker consistency, CoVoMix2 attains the best SA-WER and SA-SIM, highlighting its ability to maintain both linguistic accuracy and consistent speaker identity across multi-turn interactions.
@@ 5.2 Speaker Consistency
To evaluate speaker consistency across turns, we selected a long dialogue sequence containing 12 speech turns, with each speaker contributing 6 utterances.
Figure 4(a) shows the average speaker similarity between each generated turn and its corresponding prompt. While Sesame achieves relatively high similarity in some turns, its performance is inconsistent. This may be due to its design, which generates each utterance independently, relying on a prompt placed only at the beginning. As the dialogue progresses, the prompt information may be forgotten or get confused, especially in longer contexts. MoonCast, by contrast, demonstrates consistently low similarity. Although it incorporates prompts at the beginning of each turn, this additional conditioning does not improve speaker accuracy, indicating limited benefit despite the increased computational overhead. In comparison, CoVoMix2 achieves the highest average similarity with the lowest variance, even though the prompt is provided only once at the beginning of the dialogue. This result demonstrates the robustness in maintaining speaker identity over long conversations.
Figure 4(b) further investigates intra-dialogue speaker consistency by measuring pairwise speaker similarity across all turns from the same speaker. A more uniform and consistent color distribution along the rows and columns reflects stronger identity preservation. Both MoonCast and Sesame display noticeable variability, indicating timbre drift or identity shift during generation. In contrast, our model demonstrates consistently high pairwise speaker similarity across all turns, highlighting its effectiveness in preserving speaker timbre throughout multi-turn dialogues.
@@ 6 Ablation Studies and Extension
We conducted extensive ablation studies in Appendix C to validate our design choices across data composition, training stages, and special token representation. Results show that pretraining on monologue data is essential for learning stable pronunciation, while data mixing strategy improves generalization to diverse dialogue patterns. Simulated dialogue data may introduce noise. Additionally, we demonstrate that a brief fine-tuning stage using just 20 minutes of human-annotated data can significantly boost transcription accuracy and speaker consistency.
@@ Appendix A Model Comparison
We compare our proposed CoVoMix2 with two baseline models: MoonCast [10]555https://github.com/jzq2000/MoonCast and Sesame [8] 666https://huggingface.co/spaces/sesame/csm-1b. The detailed model comparisons are shown in Table 2, where N/A means not available.
@@ Appendix B Monologue Performance Comparison
To ensure that the dialogue generation models retain strong performance in monologue settings, we evaluate all systems on an audiobook-style monologue test set. The transcription of the audio prompts is obtained using Whisper-large-v3 [47]777https://huggingface.co/openai/whisper-large-v3. As shown in Table 3, our CoVoMix2 trained solely on monologue data, outperforms all baseline models. Notably, after the second-stage training on dialogue data, CoVoMix2 maintains its monologue generation capabilities without much degradation. In contrast, Sesame and MoonCast exhibit degraded performance due to hallucination issues, frequently producing repetitive or semantically irrelevant content.
@@
Appendix C Ablation Studies
@@
C.1 Data Mixing Strategy
@@ C.2 Training Stages
C.3 Silence Token Representation
@@
Appendix D Subjective Evaluation

### v2

@@ 1 Introduction
The field of speech synthesis has witnessed remarkable progress in recent years, particularly in zero-shot text-to-speech (TTS) systems that can generate high-quality, natural-sounding speech in voices not seen during training [1, 2, 3, 4, 5, 6, 7, 8]. These advancements have enabled applications such as personalized virtual assistants, audiobook narration, and interactive voice response systems. However, extending these capabilities to multi-talker dialogue generation, where multiple speakers engage in natural, dynamic, multi-turn conversations, remains a significant challenge.
@@ 2.3 Conversational Speech Synthesis
Conversational Speech Synthesis (CSS) focuses on generating individual utterances that are contextually appropriate within a dialogue, typically from a single speaker. Unlike full dialogue generation, CSS produces each utterance one-by-one, conditioned on previous dialogue history, rather than synthesizing an entire conversation simultaneously [30, 31, 32, 33].
For instance, GPTTalker [34] models multimodal dialogue context by converting history into discrete tokens processed by a language model to generate expressive, context-aware responses. Sesame [10] adopts two AR transformers [35]: a multimodal transformer backbone processes text and audio tokens, followed by a audio decoder transformer that reconstructs high-quality speech. Although Sesame supports dialogue-style synthesis, it fundamentally generates each speaker’s utterance sequentially given the previous dialogue context, rather than modeling simultaneous multi-speaker interaction.
@@ 3.3 Prompt-Level Random Masking
As shown in Figure 1 and 2, for each training sample, we first find all the available monologue segments for each speaker as the prompt candidates. We then randomly select a prompt segment from the candidates. The chosen prompts are concatenated at the very beginning with the masked training sample using a separator token to construct the prompt sequence . Moreover, in the text stream , we use special tokens and as indicators (not text tokens) to replace the transcription of the prompt, distinguishing which segments belong to each speaker.
@@ 3.5 Training Strategy: Curriculum Learning and Data Mixing
To improve robustness and generalization, built on the large scale of monologue dataset, we design the data mixing strategy, using various sources of data during the second stage training. Specifically, we mix data from several sources, including ASR-transcribed podcast dialogues, audiobook-style single-speaker datasets and simulated overlapped dialogues to support overlapping capability. Benefited from this data diversity, CoVoMix2 does not require any human-annotated dialogues to get satisfied and natural results. For further enhancement, we demonstrate in Appendix D.2 that fine-tuning the model with just 20 minutes of clean, human-annotated dialogue can significantly boost performance.
@@ 4.1 Training and Inference Data Preparation
To train CoVoMix2, we curated a diverse internal dataset comprising both dialogues and clean monologue data. The core training corpus consists of 3,000 hours of internal English podcast data, which we processed with 4 steps, as shown in Figure 3.
First, we applied Microsoft speech enhancement API 222https://learn.microsoft.com/en-us/azure/ai-services/speech-service/audio-processing-overview to remove background noise and music. Second, we implemented automatic speech recognition and speaker diarization. The diarization results were used to assign speaker identities to each utterance with timestamps. Specifically, we used the Deepgram API [38]333https://deepgram.com/ to obtain word-level transcriptions and speaker labels. Third, following [11]444https://github.com/vivian556123/NeurIPS2024-CoVoMix/tree/main, long dialogues were segmented into shorter clips, each involving two or one speakers, to improve data quality and training efficiency. Finally, we simulate dialogue segments using the monologue datasets. By pairing utterances from different speakers and introducing controlled overlap or silence ratios, we generate synthetic two-speaker dialogues. Detailed data-processing pipeline is open-sourced 555https://github.com/vivian556123/covomix2-dataprep.git.
@@
In order to evaluate the model performance, we design a dialogue test set 666https://github.com/vivian556123/covomix2-dialogue-testset.git, containing 1000 dialogue transcriptions from Dailydialog [41] and the acoustic prompts are from Librispeech-test-clean [42]. We also use samples from this dialogue dataset for subjective evaluation.
@@ 4.3 Baseline and Evaluation Metrics
To comprehensively assess the recognition accuracy and speaker consistency, we adopt the following objective evaluation metrics: Real-Time Factor (RTF), Word Error Rate (WER), Speaker-Aware Word Error Rate (SA-WER) [45], Speaker-Aware Speaker Similarity (SA-SIM) and UTMOS [46]. Specifically, We measure the RTF on a single NVIDIA A100 machine. We utilize Microsoft Fast Transcription API 777https://learn.microsoft.com/en-us/azure/ai-services/speech-service/speech-to-text#fast-transcription as automatic speech recognition and diarization tool to transcribe the generated speech, and we calculate the SA-WER for each word and their corresponding speaker identity. The SA-SIM enhances SIM by ensuring that the similarity is computed between embeddings attributed to the speaker identity detected by the diarization model. We utilize WavLM-TDNN [47] to extract the speaker embeddings.
We perform a human evaluation on the generated dialogue examples. We conducted a Comparable Mean Opinion Score (CMOS) experiment to assess user preference in terms of speaker turn handling, interactivity, fluency, and coherence. 15 professional linguistic experts provide judges for all subjective evaluations. They provide a rating to the second audio, which is randomly selected from a pair of audios, in the (-3 to +3) range. Detailed instructions are in Appendix E.
@@ 5.1 Objective and Subjective Evaluation Results
In terms of content accuracy and speaker consistency, CoVoMix2 attains the best SA-WER and SA-SIM, highlighting its ability to maintain linguistic accuracy and consistent speaker identity across multi-turn interactions.
@@ 5.2 Speaker Consistency
To evaluate speaker consistency across turns, we selected three long dialogue sequence containing 12 speech turns, with each speaker contributing 6 utterances.
Figure 4(a) shows the average speaker similarity between each generated turn and its corresponding prompt. While Sesame achieves relatively high similarity in some turns, its performance is inconsistent. This may be due to its design, which generates each utterance independently, relying on a prompt placed only at the beginning. As the dialogue progresses, the prompt information may be forgotten or get confused, especially in longer contexts. MoonCast, by contrast, demonstrates consistently low similarity. Although it incorporates prompts at the beginning of each turn, this additional conditioning does not improve speaker accuracy, indicating limited benefit despite the increased computational overhead.
In comparison, CoVoMix2 achieves the highest average similarity with the lowest variance, even though the prompt is provided only once at the beginning of the dialogue. This result demonstrates the robustness in maintaining speaker identity over long conversations.
Figure 4(b) further investigates intra-dialogue speaker consistency by measuring pairwise speaker similarity across all turns from the same speaker in a long dialogue. A more uniform and consistent color distribution along the rows and columns reflects stronger identity preservation. Both MoonCast and Sesame display noticeable variability, indicating timbre drift or identity shift during generation. In contrast, our model demonstrates consistently high pairwise speaker similarity across all turns, highlighting its effectiveness in preserving speaker timbre throughout multi-turn dialogues.
@@ 6 Ablation Studies and Extension
We conducted extensive ablation studies in Appendix D to validate our design choices across data composition, training stages, and special token representation. Results show that pretraining on monologue data is essential for learning stable pronunciation, while data mixing strategy improves generalization to diverse dialogue patterns. Simulated dialogue data may introduce noise.
Our experiments further demonstrate the robustness of CoVoMix2 to monologue generation, generation with very short prompts, and long dialogue generation. Specifically, it maintains stability even when other baselines fail with prompts under 10 seconds. Despite being trained on data under 30 seconds, CoVoMix2 is capable of generating stable conversations up to three times longer (approximately 90 seconds). Moreover, we demonstrate that a brief fine-tuning stage using just 20 minutes of human-annotated data can significantly boost transcription accuracy and speaker consistency.
@@ Appendix A Model Comparison
We compare our proposed CoVoMix2 with two baseline models: MoonCast [12]888https://github.com/jzq2000/MoonCast and Sesame [10] 999https://huggingface.co/spaces/sesame/csm-1b. The detailed model comparisons are shown in Table 2, where N/A means not available.
@@ Appendix B Monologue Performance Comparison
To ensure that the dialogue generation models retain strong performance in monologue settings, we evaluate all systems on an audiobook-style monologue test set. The transcription of the audio prompts is obtained using Whisper-large-v3 [51]101010https://huggingface.co/openai/whisper-large-v3. As shown in Table 3, our CoVoMix2 trained solely on monologue data, outperforms all baseline models. Notably, after the second-stage training on dialogue data, CoVoMix2 maintains its monologue generation capabilities without much degradation. In contrast, Sesame and MoonCast exhibit degraded performance due to hallucination issues, frequently producing repetitive or semantically irrelevant content.
@@
Appendix C Extended Dialogue Performance Comparison
C.1 Dialogue Performance Comparison with Prompts of different length
To demonstrate that our model maintains a stable advantage even with very long or very short prompts, we conducted an additional experiment. We selected 20 speech clips, each longer than 20 seconds, from 20 different speakers in the LibriSpeech dataset [42]. We then trimmed these clips to lengths of 3 to 18 seconds. For transcription, we used Whisper-large-v3, and for the text component, we used 100 dialogues from the Dailydialog dataset. As shown in Figure 6, our model’s SA-WER and SA-SIM performance remains consistently strong across all prompt lengths. This further confirms that our model performs reliably well, regardless of how long the input prompt is.
C.2 Dialogue Performance Comparison with Different Synthesized Length
Due to resource constraints, our training data currently consists of segments less than 30 seconds in duration. Nevertheless, our model demonstrates the ability to infer dialogues longer than 30 seconds, and our test set includes numerous dialogues exceeding one minute.
To further investigate long dialogue generation, we manually designed four dialogues with varying durations. As shown in Table 4, we observed that both CoVoMix2 and Mooncast exhibit degraded performance after 90 seconds, while Sesame fails to generate dialogues longer than 120 seconds. Additionally, Sesame showed unstable results even for dialogues under 30 seconds, primarily due to speaker confusion issues.
| System | 30s | 60s | 90s | 120s |
| Mooncast | 5.81/0.647 | 4.57/0.576 | 13.95/0.487 | 19.75/0.516 |
| Sesame | 24.41/0.677 | 5.22/0.585 | 3.98/0.263 | Failed |
| CoVoMix2 | 6.97/0.686 | 5.88/0.648 | 14.74/0.618 | 14.64/0.627 |
C.3 Dialogue Performance Comparison under Real-world Scenarios
We designed an additional test set to use more realistic conversation style speeches as prompts. The speaker prompts for this new set are derived from 10 distinct real-life dialogues from the NCSSD dataset-CEN [34]. This dataset is collected from the internet and features prompts with natural conversational prosody, including various acoustic scenarios such as background noise. The corresponding text content for this test set is sourced from the DailyDialog dataset [41].
As presented in Table 5, our model, CoVoMix2, consistently demonstrates superior performance compared to baseline models even under these challenging real-world conditions.
| System | SA-WER | SA-SIM | UTMOS |
| Mooncast | 26.26 | 0.276 | 2.09 |
| Sesame | 25.19 | 0.250 | 1.80 |
| CoVoMix2 | 9.32 | 0.322 | 2.89 |
Appendix D Ablation Studies
@@
D.1 Data Mixing Strategy
@@ D.2 Training Stages
D.3 Silence Token Representation
@@
Appendix E Subjective Evaluation